OpenStudy (anonymous):

solve the equation 4x+7=-9

OpenStudy (el_tucan):

you want to solve for x. move the 7 over to isolate 4x. So -9 - 7= -16.

OpenStudy (anonymous):

ok thanks bro

OpenStudy (anonymous):

4x - 9 = 7 4x = 7 + 9 4x = 16 x = 16/4 x = 4